2. A load of 50 kg is applied to the lower end of a steel rod 80 cm long and 0.6 cm in diameter. How much will the rod stretch? Y = 190 GPa for steel.
1
Expert's answer
2020-07-13T11:47:39-0400

The stretch of a rod is given by the following expression:


"\Delta L = \dfrac{FL}{AY}"

where "F = mg" is the force applied to the lower end of the steel rod, "L = 80cm = 0.8m" is the length of the rod, "A = \pi d^2 /4" is a cross-section area of the rod and "Y = 190\times 10^{9}Pa" is the Young modulus. Thus, obtain:

"\Delta L = \dfrac{FL}{AY} = \dfrac{4mgL}{\pi d^2Y} = \dfrac{4\cdot 50\cdot 9.8\cdot 0.8}{\pi\cdot 0.006^2\cdot 190\cdot 10^9} \approx 0.073mm"

Answer. 0.073mm.